OpenStudy (anonymous):

Derivative of: (x)(sqroot: 4-x)

OpenStudy (kc_kennylau):

\(\dfrac d{dx}x\sqrt{4-x}\) am I correct

OpenStudy (kc_kennylau):

Use product rule and then chain rule :)

OpenStudy (anonymous):

Yes, you're correct. When I found the derivative I got 4-3x over (2)(sqroot: 4-x) but the answer says that in the numerator 4 should be an 8..

OpenStudy (kc_kennylau):

\[\frac d{dx}x\sqrt{4-x}\]\[=1\cdot\sqrt{4-x}+x\cdot\frac{-1}{2\sqrt{4-x}}\]\[=\frac{2(4-x)}{2\sqrt{4-x}}-\frac x{2\sqrt{4-x}}\]

OpenStudy (kc_kennylau):

I hope that explains the \(8\) :)

OpenStudy (anonymous):

In the first step, where did the -1 come from in the numerator?

OpenStudy (kc_kennylau):

chain rule, \(\dfrac d{dx}(4-x)=-1\) :)

OpenStudy (anonymous):

Okay I see it - Thank you so much!

OpenStudy (kc_kennylau):

no problem :)